Wildcat Corporation recently disclosed the following financial​ information:

​Earnings/revenue                                                                                      

​$1,344,470

Assets                                                                                                                                               

​$7,600,000

Liabilities                                                                                                                                    

​$1,469,019

Shares outstanding                                                                           392 comma 884Market price                                                                                                                

​$30.00 per share

Calculate the​ price-to-book ratio, the​ price/earnings ratio, and the book value per share for each of the following separate​ scenarios:

a.

Based on current information

b. Earnings fall to

​$896,313

c. Liabilities increase to

​$2,798,970

d. The company does a​ three-for-one stock split with no change in market capitalization

e. The company repurchases 20 percent of the outstanding​ stock, incurring additional liability to finance the purchase.

a. Based on current​ information, the book value per share is

​$nothing.

​(Round to the nearest​ cent.)

Based on current​ information, the​ market-to-book (price/book) ratio is

nothing

. ​(Round to two decimal​ places.)

Based on current​ information, the​ price/earnings ratio is

nothing .

​(Round to one decimal​ place.)

b. If earnings fall to

​$896,313 ,

the book value per share is

​$nothing.

​(Round to the nearest​ cent.)
Answer #1

BVPS = BV of Equity / No. of Shares

Price-to-Book Ratio = Market Price of Equity / Book Value of Equity

Price/Earnings Ratio = Market Price of Equity / Earnings per Share

a). BVPS = [$7,600,000 - ​$1,469,019] / 392,884 = $15.61 per share

Price-to-Book Ratio = $30 / $15.61 = 1.92

Price/Earnings Ratio = $30 / [$1,344,470 / 392,884] = 8.77

b). BVPS = [$7,600,000 - ​$1,469,019] / 392,884 = $15.61 per share

Price-to-Book Ratio = $30 / $15.61 = 1.92

Price/Earnings Ratio = $30 / [$896,313 / 392,884] = 13.15

c). BVPS = [$7,600,000 - ​$2,798,970] / 392,884 = $12.22 per share

Price-to-Book Ratio = $30 / $12.22 = 2.45

Price/Earnings Ratio = $30 / [$1,344,470 / 392,884] = 8.77

d). BVPS = [$7,600,000 - ​$1,469,019] / [392,884 x 3] = $5.20 per share

Price-to-Book Ratio = $30 / $5.20 = 5.77

Price/Earnings Ratio = $30 / [$1,344,470 / (392,884 x 3)] = 26.29

e). BVPS = [$7,600,000 - $1,469,019 - {$30 x (392,884 x 0.20)}] / [392,884 - (20% x 392,884)]

= $3,773,677 / 314,307.20 = $12.01 per share

Price-to-Book Ratio = $30 / $12.01 = 2.50

Price/Earnings Ratio = $30 / [$1,344,470 / {392,884 x (1 - 0.20)}] = $30 / $4.28 = 7.01